# Instant 3D Human Avatar Generation using Image Diffusion Models

### v1

@@ Abstract
We present AvatarPopUp, a method for fast, high quality 3D human avatar generation from different input modalities, such as images and text prompts and with control over the generated pose and shape. The common theme is the use of diffusion-based image generation networks that are specialized for each particular task, followed by a 3D lifting network. We purposefully decouple the generation from the 3D modeling which allow us to leverage powerful image synthesis priors, trained on billions of text-image pairs. We fine-tune latent diffusion networks with additional image conditioning to solve tasks such as image generation and back-view prediction, and to support qualitatively different multiple 3D hypotheses. Our partial fine-tuning approach allows to adapt the networks for each task without inducing catastrophic forgetting. In our experiments, we demonstrate that our method produces accurate, high-quality 3D avatars with diverse appearance that respect the multimodal text, image, and body control signals. Our approach can produce a 3D model in as few as 2 seconds, a four orders of magnitude speedup w.r.t. the vast majority of existing methods, most of which solve only a subset of our tasks, and with fewer controls, thus enabling applications that require the controlled 3D generation of human avatars at scale. The project website can be found at https://www.nikoskolot.com/avatarpopup/.
@@ 1 Introduction
Our proposed decoupling strategy allows us to maximally exploit available data sources with different levels of supervision. We first fine-tune a pretrained Latent Diffusion network to generate images of people based on textual descriptions and with additional control over the desired pose and shape. This step does not require any ground truth 3D data for supervision and enables scaling our image generator to web scale data of images of people in various poses. Next, we leverage a small-scale dataset of scanned 3D human assets and fine-tune a second latent diffusion network to learn the distribution of back side views conditioned on a front view image of the person. We optionally also condition on a textual description that can naturally complement the evidence available in the front view image. Furthermore, we propose a novel fine-tuning strategy that prevents overfitting to the new datasets. Finally, we design and train a 3D reconstruction network that predicts a textured 3D shape in the form of an implicit signed distance field given the pair of front and back views and optionally 3D body signals. The resulting cascaded method, which we call AvatarPopUp supports a wide range of 3D generation and reconstruction tasks: First, ig enables fast and interactive 3D generation of assets at scale, see Fig. 1. Second, we can repurpose parts of the cascade for image-based 3D reconstruction at state-of-the-art quality. Finally, we demonstrate how AvatarPopUp can be used for creative editing tasks exemplified in 3D virtual try-on with body shape preservation. To summarize, our main contributions are:
@@ 3.2 Back View Generation
One could try to lift the generated front views from the previous stage to 3D directly by applying a single-image 3D reconstruction method like PHORHUM [8]. However, because of the inherent ambiguity of the problem this will result in significant loss of geometric detail and blurry textures for the non-visible body surfaces. To avoid it, we propose to fine-tune again a latent diffusion network with the same strategy as in the previous section. This time the additional image conditioning is a front view and optionally a text prompt, and we train the network to learn the distribution of back views conditioned on the front view. The additional text prompt can be used in cases where it is desired to additionally guide the generation by very specific properties. Fig. 3 shows different back sides sampled from the conditional distribution. We also show that the additional text inputs are useful in modulating certain parts of the generation that are not immediately deducible from the front image, such as hairstyles or specific patterns.
@@
+ “gray hair
+ “with a tattoo
@@ 4.1 Text-to-3D Generation
In Sec. 4.1 we use CLIP to evaluate our model against other text-to-3D generation methods. In general, CLIP-based metrics are not indicative of the generated image quality, because they only consider the alignment with the text, and often over-saturated images with extreme details tend to have high CLIP scores. To further demonstrate that our method generates higher quality avatars, we also include a qualitative comparison in Fig. 5.
@@
– – 0.03 0.08 3m CHUPA [31]
0.58 0.73 0.08 0.17
$\sim$2s Ours
0.62 0.77
0.11 0.17 $\sim$10s Ours (high quality)
@@
Figure 5: Comparisons with text-to-3d human generation methods. Our method generates high quality results that respect the text prompt well, at a fraction of the others’ runtime, cf. Sec. 4.2. TADA’s results appear unnatural; DreamHuman failed for one subject and produces oversaturated colors; CHUPA failed to respect the prompt.
@@ 4.2 Single-image 3D Reconstruction
While not specifically designed for 3D reconstruction, our method achieves state-of-the-art performance also for this task. The evaluation setup is the following: given an input image $I$, we draw one random sample from the back view image generator network, and then feed the pair of front/back images to our 3D reconstruction network. For all methods we extract a textured 3D mesh using Marching Cubes and report numerical results in Sec. 4.2. Furthermore, we show qualitative results in Fig. 6. Notably, our method not only performs on par numerically and qualitatively on reconstructed front views, but also generates highly detailed back view texture and geometry. Finally, we also compare with the optimization-based method TeCH [27] in Fig. 7. TeCH produces detailed front and back geometry but also also exhibits problems at times, rooted in its 3D pose estimation method. Most importantly, TeCH runs for several hours per instance, while ours computes results in a single feed-forward pass, in only a few seconds.
{NiceTabular}
[…]
@@
7.1 0.50 0.72 – – 29.6 57.2 PIFu [60]
6.2 0.52 0.74 – – 26.9 52.5 PIFuHD [61]
7.0 0.48 0.72 – – 36.1 49.0 ICON [71]
3.4 0.58 0.76 – – 18.3 36.1 ECON [70]
@@
0.91 0.70 0.86
11.0 20.0 12.8 25.8 Ours
Table 3: Numerical comparisons with single-view 3D reconstructions methods and ablations of our method. We mark the best and second best results. All Chamfer metrics are ×$\times 10^{-3}$. Not all methods generate colors. For fair comparisons, we retrained PHORHUM [8] using the same data as AvatarPopUp. We observe comparable results in terms of 3D metrics, however ours performs better on generating more realistic and diverse back views and back normals. While we used here only a single image as input, we are also able to optionally condition on text, body shape, and pose.
@@
An immediate application of our method is the option to perform 3D garment edits for a given identity. Given an input image, we first recover 3D pose and shape parameters of the person in the image. Using a text prompt and the identity preservation strategy introduced below, we can generate updated images of the same person wearing different garments or accessories. To preserve the identity of the person, we first locate the person’s head in the source image and use then Repaint [46] to out-paint a novel body for the given head while still conditioning on the estimated pose and shape parameters. We this strategy we ensure that our method not only preserves the facial characteristics but also the overall body proportions. In Fig. 9 we illustrate such editing examples. The generated 3D edits present garment details like wrinkles on both front and back views, and preserve the subjects’ facial appearance. Also, note that body shape and identity are well preserved for the subjects, even though only one image is given. While there has been a significant amount of 2D virtual try-on research [20, 86, 35], our methodology can generate consistent and highly detailed 3D meshes that can be animated and rendered from other viewpoints.
@@ 4.4 Animation
Figure 9: Identity preserving 3D avatar editing. Our method allows for editing the clothing, while preserving the identity of the generated person. In each row we’re using the image on the left as input, recover the person’s pose and shape parameters and then run our method on updated prompts using the same pose and shape conditioning. The generated avatars share the same identity (same facial appearance, same body shape, and pose) with the input image. The garments are different, have good quality textures, and present fine wrinkle detail.
@@ 4.5 Ablation Study
We compare the effectiveness of our additional pose and shape encoding inputs $\mathcal{G}$ to our reconstruction network during generation, where we have control over the target pose. To do so, we use the same set of 100 text prompts as before and for each text prompt sample a random pose and shape configuration. For each $(\tau,\boldsymbol{\theta},\boldsymbol{\beta})$ triplet we run inferences and compute 2 meshes: one using only the front and back images, and another one additionally using the dense GHUM encodings. We then evaluate the Chamfer distance between each mesh and the corresponding GHUM mesh. The model using the additional GHUM signals has an average Chamfer distance of $d_{\mathrm{with}}=1.4$ whereas the one without $d_{\mathrm{without}}=8.6$, thus validating our design choice. Not only is the control respected well, but this also allows for animation as discussed previously.
@@ A Roman marble statue
Figure 10: Partial vs. complete fine-tuning strategy. For each text prompt the first example is generated using our partial fine-tuning strategy whereas the second by fine-tuning the entire network. A partial fine-tuning has empirically proven to be more resilient to catastrophic forgetting, generating more diverse and better text aligned images, while at the same time maintaining consistency with the conditioning pose and shape.
@@ 5 Discussion & Conclusions
We present a generative tool to create 3D human assets, thus reducing the risk of scanning and using real humans for training large scale 3D generative models. AvatarPopUp generates diverse results and can be controlled with body shape, clothing and other properties, which we hope will lead to a better coverage of subject distributions and scale the amount of data available to the community.

### v2

@@ Abstract
We present AvatarPopUp, a method for fast, high quality 3D human avatar generation from different input modalities, such as images and text prompts and with control over the generated pose and shape. The common theme is the use of diffusion-based image generation networks that are specialized for each particular task, followed by a 3D lifting network. We purposefully decouple the generation from the 3D modeling which allow us to leverage powerful image synthesis priors, trained on billions of text-image pairs. We fine-tune latent diffusion networks with additional image conditioning for image generation and back-view prediction, and to support qualitatively different multiple 3D hypotheses. Our partial fine-tuning approach allows to adapt the networks for each task without inducing catastrophic forgetting. In our experiments, we demonstrate that our method produces accurate, high-quality 3D avatars with diverse appearance that respect the multimodal text, image, and body control signals. Our approach can produce a 3D model in as few as 2 seconds, a four orders of magnitude speedup w.r.t. the vast majority of existing methods, most of which solve only a subset of our tasks, and with fewer controls. AvatarPopUp enables applications that require the controlled 3D generation of human avatars at scale. The project website can be found at https://www.nikoskolot.com/avatarpopup/.
@@ 1 Introduction
Our proposed decoupling strategy allows us to maximally exploit available data sources with different levels of supervision. We first fine-tune a pretrained Latent Diffusion network to generate images of people based on textual descriptions and with additional control over the desired pose and shape. This step does not require any ground truth 3D data for supervision and enables scaling our image generator to web scale data of images of people in various poses. Next, we leverage a small-scale dataset of scanned 3D human assets and fine-tune a second latent diffusion network to learn the distribution of back side views conditioned on a front view image of the person. We optionally also condition on a textual description that can naturally complement the evidence available in the front view image. Furthermore, we propose a novel fine-tuning strategy that prevents overfitting to the new datasets. Finally, we design and train a 3D reconstruction network that predicts a textured 3D shape in the form of an implicit signed distance field given the pair of front and back views and optionally 3D body signals. The resulting cascaded method, which we call AvatarPopUp supports a wide range of 3D generation and reconstruction tasks: First, it enables fast and interactive 3D generation of assets at scale, see Fig. 1. Second, we can repurpose parts of the cascade for image-based 3D reconstruction at state-of-the-art quality. Finally, we demonstrate how AvatarPopUp can be used for creative editing tasks exemplified in 3D virtual try-on with body shape preservation. To summarize, our main contributions are:
@@ 3.2 Back View Generation
One could try to lift the generated front views from the previous stage to 3D directly by applying a single-image 3D reconstruction method like PHORHUM [8]. However, because of the inherent ambiguity of the problem this will result in significant loss of geometric detail and blurry textures for the non-visible body surfaces. To avoid this, we propose to fine-tune again a latent diffusion network with the same strategy as in the previous section. This time the additional image conditioning is a front view and optionally a text prompt, and we train the network to learn the distribution of back views conditioned on the front view. The additional text prompt can be used in cases where it is desired to additionally guide the generation by very specific properties. Fig. 3 shows different back sides sampled from the conditional distribution. We also show that the additional text inputs are useful in modulating certain parts of the generation that are not immediately deducible from the front image, such as hairstyles or specific patterns.
@@
+ “gray hair”
+ “with a tattoo”
@@ 4.1 Text-to-3D Generation
In Tab. 2 we use CLIP to evaluate our model against other text-to-3D generation methods. In general, CLIP-based metrics are not indicative of the generated image quality, because they only consider the alignment with the text, and often over-saturated images with extreme details tend to have high CLIP scores. To further demonstrate that our method generates higher quality avatars, we also include a qualitative comparison in Fig. 5.
@@
0.03 0.08 3m CHUPA [31]
0.58 0.73 0.08 0.17 $\sim$2s Ours
0.62 0.77 0.11 0.17 $\sim$10s Ours (high quality)
@@
Figure 5: Comparisons with text-to-3d human generation methods. Our method generates high quality results that respect the text prompt well, at a fraction of the others’ runtime, cf. Tab. 3. TADA’s results appear unnatural; DreamHuman failed for one subject and produces oversaturated colors; CHUPA failed to respect the prompt.
@@ 4.2 Single-image 3D Reconstruction
While not specifically designed for 3D reconstruction, our method achieves state-of-the-art performance also for this task. The evaluation setup is the following: given an input image $I$, we draw one random sample from the back view image generator network, and then feed the pair of front/back images to our 3D reconstruction network. For all methods we extract a textured 3D mesh using Marching Cubes and report numerical results in Tab. 3. Furthermore, we show qualitative results in Fig. 6. Notably, our method not only performs on par numerically and qualitatively on reconstructed front views, but also generates highly detailed back view texture and geometry. Finally, we also compare with the optimization-based method TeCH [27] in Fig. 7. TeCH produces detailed front and back geometry but also also exhibits problems at times, rooted in its 3D pose estimation method. Most importantly, TeCH runs for several hours per instance, while ours computes results in a single feed-forward pass, in only a few seconds.
3D Metrics FID $\downarrow$
@@
7.1 0.50 0.72 29.6 57.2 PIFu [60]
6.2 0.52 0.74 26.9 52.5 PIFuHD [61]
7.0 0.48 0.72 36.1 49.0 ICON [71]
3.4 0.58 0.76 18.3 36.1 ECON [70]
@@
0.91 0.70 0.86 11.0 20.0 12.8 25.8 Ours
Table 3: Numerical comparisons with single-view 3D reconstructions methods and ablations of our method. We mark the best and second best results. All Chamfer metrics are ×$\times 10^{-3}$. Not all methods generate colors. For fair comparisons, we retrained PHORHUM [8] using the same data as AvatarPopUp. We observe comparable results in terms of 3D metrics, however ours performs better on generating more realistic and diverse back views and back normals.
@@
An immediate application of our method is the option to perform 3D garment edits for a given identity. Given an input image, we first recover 3D pose and shape parameters of the person in the image. Using a text prompt and the identity preservation strategy introduced below, we can generate updated images of the same person wearing different garments or accessories. To preserve the identity of the person, we first locate the person’s head in the source image and use then Repaint [46] to out-paint a novel body for the given head while still conditioning on the estimated pose and shape parameters. With this strategy we ensure that our method not only preserves the facial characteristics but also the overall body proportions. In Fig. 9 we illustrate such editing examples. The generated 3D edits present garment details like wrinkles on both front and back views, and preserve the subjects’ facial appearance. Also, note that body shape and identity are well preserved for the subjects, even though only one image is given. While there has been a significant amount of 2D virtual try-on research [20, 86, 35], our methodology can generate consistent and highly detailed 3D meshes that can be animated and rendered from other viewpoints.
@@ 4.4 Animation
Figure 9: Identity preserving 3D avatar editing. Our method allows for editing the clothing, while preserving the identity of the generated person. In each row we’re using the image on the left as input, recover the person’s pose and shape parameters and then run our method using the shown prompts. The generated avatars share the same identity and respect the target prompts.
@@ 4.5 Ablation Study
We compare the effectiveness of our additional pose and shape encoding inputs $\mathcal{G}$ to our reconstruction network.To do so, we use the same set of 100 text prompts as before and for each text prompt sample a random pose and shape configuration. For each $(\tau,\boldsymbol{\theta},\boldsymbol{\beta})$ triplet we run inferences and compute 2 meshes: one using only the front and back images, and another one additionally using the dense GHUM encodings. We then evaluate the Chamfer distance between each mesh and the corresponding GHUM mesh. The model using the additional GHUM signals has an average Chamfer distance of $d_{\mathrm{with}}=1.4$ whereas the one without $d_{\mathrm{without}}=8.6$, thus validating our design choice. Not only is the control respected well, but this also allows for animation as discussed previously.
@@ A Roman marble statue
Figure 10: Partial vs. complete fine-tuning strategy. For each text prompt the first example is generated using our partial fine-tuning strategy whereas the second by fine-tuning the entire network. Empirically, partial fine-tuning is more resilient to catastrophic forgetting, generating more diverse and better text-aligned images.
@@ 5 Discussion & Conclusions
We present a generative tool to create 3D human assets, thus reducing the risk of scanning and using real humans for training large scale 3D generative models. AvatarPopUp generates diverse results and can lead to a better coverage of subject distributions.
@@
Supplementary Material
The Supplementary Material contains additional implementation details and experiments that were not included in the main paper due to space constraints. Additional results can be found on our project website:
https://www.nikoskolot.com/avatarpopup/.
Appendix 0.A Implementation Details
0.A.1 Latent Diffusion models
Our base text-to-image generation model that we finetune is a reimplementation of Stable Diffusion [56] with 800 million parameters trained on internal data sources. The latent space has dimensions $64\times 64\times 8$ and the input and output images are $512\times 512\times 3$. To enable image conditioning in the models, we pass the conditioning image through the latent encoder, and then we concatenate the conditioning latent with the noisy image latent $z_{t}$ at the input layer. The weights of the input layer are padded with extra channels initialized with zeros to account for the additional 8 input channels from the conditioning. We train only the parameters of the convolutional layers of the encoder. We finetune the models with the Adam [adam] optimizer using a batch size of 64 images and a learning rate of $5\times 10^{-5}$ on 16 40GB A100 GPUs, for a total of 40000 training iterations. Finetuning takes around 6 hours. We use both image and text guidance in the style of InstructPix2Pix [brooks2022instructpix2pix]. We use an text guidance weight of 7.5 and an image guidance weight of 2.0. To enable the use of classifier-free guidance, at training time we randomly drop the conditionings. We mask the input text, input image, or input image and text in a mutual exclusive way, with probability $0.05$ each.
0.A.2 3D Reconstruction model
Our network architecture is similar to PHORHUM [8]. The only modification in the encoder the number of input channels, which is $6$ in the case of the front-back model and $12$ for the model with the additional GHUM conditioning $\mathcal{G}$. Additionally, because of the ambiguity of lighting estimation, we drop the shading-albedo decomposition and output the shaded color directly. We train our model for $500$K iterations with the Adam optimizer using a batch size of 32 and a learning rate of $10^{-4}$. Training the model takes 42 hours on 16 40GB A100 GPUs. We use a subset of the original PHORHUM losses for training. We keep the on-surface loss $\mathcal{L}_{g}$, inside-outside loss $\mathcal{L}_{l}$, eikonal loss $\mathcal{L}_{e}$, and color losses $\mathcal{L}_{a}$ where we replace the albedo $\boldsymbol{a}$ with the shaded color $\boldsymbol{c}$. We omit the rendering losses $\mathcal{L}_{r}$ as in our setting we did not find them to be useful and also made training slower.
Appendix 0.B Generation Diversity
In this section we evaluate the diversity of the generations for the different text-to-3D generation methods. We use the same set of 100 generations used in the main paper. As a proxy for diversity we measure the face similarity of the generated subjects. To quantify the face similarity we use the FaceNet embeddings [schroff2015facenet]. More specifically, we detect and crop the head regions and then use FaceNet to compute the face embeddings. Given images $I_{i}$ and $I_{j}$ with embeddings $\mathbf{e}_{i}$, $\mathbf{e}_{j}$ respectively, their pairwise similarity is defined as $s_{ij}=\mathbf{e}_{i}^{T}\mathbf{e}_{j}\in[0,1]$. The average pairwise similarity $\mathcal{S}$ over a set of images $\mathcal{D}$ is then defined as:
$\mathcal{S}=\frac{\sum_{I_{i}\in\mathcal{D}}\sum_{I_{j}\in\mathcal{D}\backslash\{I_{i}\}}s_{ij}}{|\mathcal{D}|\cdot(|\mathcal{D}|-1)}.$ (4)
Intuitively, a high value of $\mathcal{S}$ means that the generated faces are similar to each other. We also consider the maximum pairwise similarity between an image $I_{i}$ and a reference dataset $\mathcal{D}$ defined as $s_{i}=\max_{I_{j}\in\mathcal{D}}s_{ij}$.
The similarity metric $s_{i}$ quantifies whether the face in image $I_{i}$ is similar to some face from dataset $\mathcal{D}$.
As shown on Table 0.B our method is able to generate more diverse faces than representative optimization-based methods like DreamHuman [32] or TADA[37]. We use the similarity between 100 randomly selected training subjects as reference, and our method is able to generate people with comparable similarity scores.
Additionally, we test whether our method overfits on training identities, by comparing the average maximum similarities of our generations with $200$ randomly sampled 3D models from the training and test sets. As reported on Table 0.B, our generated avatars have the same pairwise similarity scores with either models, thus showing that our model did not overfit on the training set identities.
{NiceTabular}
c|l[colortbl-like] Average pairwise face similarity $\mathcal{S}$downarrow &
0.64 DreamHuman [32]
0.78 TADA [37]
0.62 Ours
0.55 Random training subjects (lower bound)
Table 4: Generation diversity evaluation. We mark the best and second best results. Our method is able to generate faces with larger diversity than the baselines. For reference, we also report the average face similarity for training subjects.
{NiceTabular}
cc|l[colortbl-like] Average Maximum Similarity Median Maximum Similarity
0.67 0.67 Train subjects
0.67 0.67 Test subjects
0.88 0.81 Generations (self-similarity)
Table 5: Evaluating training set memorization. We evaluate the similarity of our generated faces with those from the training and test sets. The results show that our model did not overfit on the training set identities.
Appendix 0.C Additional Qualitative Results
In this section we show additional qualitative results that we could not include in the main paper due to space constraints.
0.C.1 Relightable Avatar Generation
By design, our image generation models produce images of people with shading. We additionally experiment with generating albedo images instead of shaded ones, and in this way we can create 3D avatars that can then be relighted in different environments. We teach our model to produce albedo images by randomly substituting the shaded model renderings with unshaded ones at training time, and also appending “uniform lighting” to the text prompt. In Figure 12 we show example generations that are rendered in different HDRI environments.
0.C.2 Semantic editing
We show additional results for the task of semantic editing. We explore 2 scenarios: changing only specific garments on the body while preserving the rest of the appearance and changing the identity of the person wearing the outfit. The input to our method is an image of a person, editing instructions and corresponding semantic segmentation masks. The results are shown on Figure 11
Figure 11: Semantic editing. In each row we’re using the image on the left as input, recover the person’s pose and shape parameters and then run our method on updated prompts using the same pose and shape conditioning. In the first row we use additional clothing segmentation masks to perform edits only in specific body regions. In the second row we mask out the person and then generate new people in the same pose wearing the same outfit.
Figure 12: Albedo generation and relighting. The first row shows 8 avatars generated by probing our method to generate albedo instead of shaded colors. The next 3 rows shows the results of rendering the avatars in different HDRI environments.